OpenStudy (anonymous):

How do you solve this: log base of 2 (x+4) = 3 - log base of 2 (x-2) ?

OpenStudy (anonymous):

\[\log_{2} (x+4) = 3 - \log_{2} (x-2)\]

OpenStudy (anonymous):

log(x + 4) / log(2) = 3 - log(x - 2) / log(2)

OpenStudy (anonymous):

Ok, but now what?

OpenStudy (anonymous):

Do you want the solution?

OpenStudy (anonymous):

or do you want the alternate form assuming x is positive?

OpenStudy (anonymous):

the alternate form assuming x is positive is log ((x - 2) (x + 4)) = log(8)

OpenStudy (anonymous):

after this (x-2)(x+4)=8 and simply solve for x usinq quadratic formula